TORSION-ROTATION ANALYSIS OF TORSIONAL COMBINATION BANDS BUILT ON THE METHANOL OH STRETCH OVERTONE: 2ν_{1}, 2ν_{1}+ν_{12}, AND 2ν_{1}+2ν_{12}

DAVID RUEDA, O. V. BOYARKIN, AND T. R. RIZZO, INSTITUT DE CHIMIE PHYSIQUE MOLECULAIRE, EPFL, CH - 1015 LAUSANNE; ANDREI CHIROKOLAVA, INDRANATH MUKHOPADHYAY, AND \underline{DAVID S. PERRY}, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF AKRON, AKRON, OHIO 44325.

The spectrum of the first overtone OH stretch band (2ν_{1}) in jet-cooled methanol has been measured using Infrared Laser Assisted Photofragment Spectroscopy (IRLAPS) and detailed assignments have been made. In addition to the torsional fundamental band (δv_{t}=0) it was possible to observe and assign the torsional combination band with (δv_{t}=1 and 2). These transitions are very week but can be clearly resolved in the spectra. In total, 131 transitions reaching 19 different K levels in the 2ν_{1} state have been fit to a global torsion-rotation Hamiltonian and the leading torsion-rotation parameters have been determined. Assignments for the much weaker and irregular torsional band structure for 3ν_{1}+ν_{12} and 3ν_{1}+2ν_{12} are in progress.